xx GRHayLHD

GRHayLHD

Samuel Cupp Leonardo Rosa Werneck Terrence Pierre Jacques

Zachariah Etienne

Date

Abstract

Provides a general relativistic hydrodynamic (GRHD) evolution code using GRHayLvia the GRHayLibthorn. This is a trimmed-down version of the IllinoisGRMHDthorn where magnetic fields are set to zero.

1 Introduction

This thorn provides a GRHD evolution code built on the General Relativistic Hydrodynamics Library (GRHayL), which is included in the Einstein Toolkit via the GRHayLib thorn. The core library inherits or adapts code from the original IllinoisGRMHD thorn into a modular, infrastructure-agnostic library. While the IllinoisGRMHD thorn can be set to have zero magnetic fields, they will still be evolved. This is much slower than not evolving them, and GRHayLHD provides an implementation for evolving systems without magnetic fields.

This thorn provides support for hybrid and tabulated EOS, both with and without entropy. The core functions all come from GRHayLib, ensuring that any improvements or updates to the library can be easily adopted by the thorn. The conservative-to-primitive solver is selected at runtime, allowing for users to take advantage of any Con2Prim methods provided by GRHayL.

2 Parameters

Most of the behavior of this thorn is controlled by GRHayLib, the thorn that provides GRHayL functionality within the Einstein Toolkit. However, there are several parameters in the thorn which control diagnostic or debugging features. Most are self-explanatory, but some parameters deserve a more detailed explanation.

First, many analysis or diagnostic thorns use HydroBase variables. By default, GRHayLHD never copies the data back from GRHayLHD variables to HydroBase variables, which would prevent the usage of these thorns. The ConvertxtoxHydroBasexevery parameter sets how often to copy back this data. Naturally, copying more frequently will slow down the simulation, so this should be set on the same frequency as the analysis or IO thorn that is using the HydroBase variable. We reduced the amount of data copying in GRHayLHD by using the HydroBase variables directly, which removes the need to copy them back and removes extra grid functions from memory. We still use distinct velocity variables because we use a different velocity than is defined in

HydroBase.

Next, there are several parameters for debugging or other testing. The perturbing parameters give users the ability to affect either the initial data or evolution data at runtime without slowing down standard simulations. The perturbxinitialxdata parameter controls whether the initial data from HydroBase is perturbed, and the perturbxeveryxcon2prim parameter controls whether the conservative variables are perturbed before every con2prim call. The magnitude of these perturbations is set by randomxpert.

3 Parameters

11 convertxtoxhydrobasexevery 10 Scope: private01  INT

11 Description: How often to convert GRHayLHD primitive variables to HydroBase (Valencia formulation) primitive variables? Needed for some ET-based diagnostics. NOT needed for pure GRHayLHD runs.

11  Range  10              01  Default: (none)
11 0:*  10 zero (disable) or positive (every N iterations)

11 matterxbc  10 Scope: private01  KEYWORD

11 Description: Chosen Matter boundary condition

11 Range 10 01 Default: outflow
11 copy  10 Copy BCs
11 outflow  10 Copy BCs with additional outflow BC for velocities
11 frozen  10 Frozen boundaries

11

11 perturbxeveryxcon2prim 10 Scope: private01  BOOLEAN

11 Description: Random perturbation applied every time con2prim runs

11                       10              01  Default: no

11

11 perturbxinitialxdata  10 Scope: private01  BOOLEAN

11 Description: Random perturbation to initial data

11                       10              01  Default: no

11

11 randomxpert  10 Scope: private01  REAL

11 Description: Magnitude of random perturbations

11 Range 10 01 Default: (none)
11 *:*  10 Anything goes.

11

11 randomxseed  10 Scope: private01  INT

11 Description: Random seed for random, generally roundoff-level perturbation on initial data. Seeds srand(), and rand() is used for the RNG.

11  Range  10              01  Default: (none)
11 0:99999999  10 Anything unsigned goes.

11

11 symmetry  10 Scope: private01  KEYWORD

11 Description: Currently only no symmetry supported, though work has begun in adding equatorial-symmetry support.

11  Range  10              01  Default: none
11 none  10 no symmetry, full 3d domain

11

11 updatextmunu  10 Scope: private01  BOOLEAN

11 Description: Update Tmunu, for RHS of Einstein’s equations?

11                       10              01  Default: yes

11

11 verbose  10 Scope: private01  KEYWORD

11 Description: Determines how much evolution information is output

11 Range 10 01 Default: yes
11 no  10 Complete silence
11 yes  10 ”Outputs information about conservative-to-prim itive fixes, failures, etc.”

11

11 eosxtype  10 Scope: shared from GRHAYLIB01KEYWORD

11 evolvexentropy  10 Scope: shared from GRHAYLIB01BOOLEAN

11

4 Interfaces

General Implements: grhaylhd

Inherits: admbase tmunubase hydrobase grhaylib

Grid Variables

4.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  grmhdxvelocities  10                    01  compact 11 0
11                       10  vx  01 description 11 Components of primitive three velocity vx. Note that vx is defined in terms of 4-velocity as: vx = ux/uˆ0 . Note that this definition differs from the Valencia formalism.
11                       10  vy  01 dimensions 11 3
11                       10  vz  01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11 u0 10 u0 01 compact 11 0
11                       10                    01 description 11 4-velocity component uˆ0
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none” checkpoint=”no”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11 grmhdxconservatives 10                    01  compact 11 0
11                       10  rhoxstar  01 description 11 Evolved hydrodynamic variables
11                       10  tau  01 dimensions 11 3
11                       10  Stildex  01 distribution 11 DEFAULT
11                       10  Stildey  01 group type 11 GF
11                       10  Stildez  01  tags 11 prolongation=”ENO”
11                       10                    01 timelevels 11 3
11                       10                    01 variable type 11 REAL
11 entxstar 10 entxstar 01 compact 11 0
11                       10                    01 description 11 Evolved entropy variable
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 prolongation=”ENO”
11                       10                    01 timelevels 11 3
11                       10                    01 variable type 11 REAL
11 yexstar 10 Yexstar 01 compact 11 0
11                       10                    01 description 11 Evolved electron fraction variable
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 prolongation=”ENO”
11                       10                    01 timelevels 11 3
11                       10                    01 variable type 11 REAL
11 grmhdxconservativesxrhs 10                    01  compact 11 0
11                       10  rhoxstarxrhs  01 description 11 Right-hand side for grmhdxconservatives variables
11                       10  tauxrhs  01 dimensions 11 3
11                       10  Stildexxrhs  01 distribution 11 DEFAULT
11                       10  Stildeyxrhs  01 group type 11 GF
11                       10  Stildezxrhs  01  tags 11 prolongation=”none” Checkpoint=”no”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 entxstarxrhs 10 entxstarxrhs 01 compact 11 0
11                    10                    01 description 11 Right-hand side for entxstar
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 yexstarxrhs 10 Yexstarxrhs 01 compact 11 0
11                    10                    01 description 11 Right-hand side for Yexstar
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 grmhdxfluxxtemps 10                    01  compact 11 0
11                    10  rhoxstarxflux 01 description 11 Flux variables for computing RHSs
11                    10  tauxflux  01 dimensions 11 3
11                    10  Stildexxflux  01 distribution 11 DEFAULT
11                    10  Stildeyxflux  01 group type 11 GF
11                    10  Stildezxflux  01  tags 11 prolongation=”none” Checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 entxstarxflux 10 entxstarxflux 01 compact 11 0
11                    10                    01 description 11 Flux variable for entxstar
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 yexstarxflux 10 Yexstarxflux 01 compact 11 0
11                    10                    01 description 11 Flux variable for Yexstar
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 failurexchecker 10 failurexchecker 01 compact 11 0
11                    10                    01 description 11 Gridfunction to track conservative-to-primitives solver fixes. Beware that this gridfunction is overwritten at each RK substep.
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”none” Checkpoint=”no” InterpNumTimelevels=1
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

Uses header: Symmetry.h GRHayLib.h

5 Schedule This section lists all the variables which are assigned storage by thorn GRHayLET/GRHayLHD. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 Conditional:
11 grmhdxconservatives[3] grmhdxvelocities 11 Yexstar[3] Yexstarxrhs Yexstarxflux
11 u0 grmhdxconservativesxrhs  11 entxstar[3] entxstarxrhs entxstarxflux
11 grmhdxfluxxtemps failurexchecker  11 x

11 x  11 x

11

Scheduled Functions

MoLxRegister (conditional) x grhaylhdxregistervars x register evolved, rhs variables in grhaylhd for mol x

11 x 11 Language: 11 c
11 x 11 Options: 11 meta
11 x 11 Type:  11 function

11

CCTKxBASEGRID (conditional) x grhaylhdxinitsymbound x schedule symmetries x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

GRHayLHDxCon2Prim (conditional) x grhaylhdxperturbxconservatives x perturb conservatives before con2prim x

11 x 11 After: 11 grhaylhdxsyncxconservatives
11 x 11 Before: 11 grhaylhdxconservsxtoxprims
11 x 11 Type: 11 group

11

GRHayLHDxCon2Prim (conditional) x grhaylhdxconservsxtoxprims x compute primitive variables from conservatives x

11 x 11 After: 11 grhaylhdxsyncxconservatives
11 x 11 Type: 11 group

GRHayLHDxCon2Prim (conditional) x grhaylhdxouterxboundaries x apply selected boundary conditions to the primitive variables x

11 x 11 After: 11 grhaylhdxconservsxtoxprims
11 x 11 Type: 11 group

11

AddToTmunu (conditional) x grhaylhdxcomputextmunu x compute stress-energy tensor x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 grmhdxvelocities

11 x 11 x  11 u0

11 x 11 Type:  11 function
11 x 11 Writes: 11 tmunubase::stressxenergyxscalar(everywhere)
11 x 11 x  11 tmunubase::stressxenergyxvector(everywhere)
11 x 11 x  11 tmunubase::stressxenergyxtensor(everywhere)

11

MoLxCalcRHS (conditional) x grhaylhdxrhs x evaluate rhss grhd equations x

11 x 11 Type: 11 group

11

GRHayLHDxRHS (conditional) x grhaylhdxevaluatexsourcesxrhs x evaluate source terms in grhd rhss x

11 x 11 Type: 11 group

11

GRHayLHDxRHS (conditional) x grhaylhdxevaluatexfluxesxrhs x evaluate flux terms in grhd rhss x

11 x 11 After: 11 grhaylhdxevaluatexsourcesxrhs
11 x 11 Type: 11 group

GRHayLHDxRHS (conditional) x convertxgrhaylhdxtoxhydrobase x convert needed hydrobase variables for nrpyleakage x

11 x 11 After:  11 grhaylhdxevaluatexfluxesxrhs
11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::vel(everywhere)

11

CCTKxANALYSIS (conditional) x convertxgrhaylhdxtoxhydrobase x convert grhaylhd-native variables to hydrobase x

11 x 11 After:  11 mlxbssnxevolcalcgroup
11 x 11 Before: 11 computexbixb2xpoynxfluxet
11 x 11 x  11 convertxtoxmhdx3velocity
11 x 11 x  11 particlextraceret
11 x 11 x  11 volumeintegralgroup
11 x 11 Language: 11 c
11 x 11 Options: 11 global-early
11 x 11 x  11 loop-local
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::vel(everywhere)
11 x 11 x  11 hydrobase::wxlorentz(everywhere)

11

GRHayLHDxprimsxtoxconservs (conditional) x grhaylhdxhybridxentropyxprimsxtoxconservs x entropy+hybrid version of grhaylhdxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)

HydroBasexPrim2ConInitial (conditional) x grhaylhdxprim2con2prim x convert hydrobase initial data to grhaylhd variables x

11 x 11 Type: 11 group

11

GRHayLHDxconservsxtoxprims (conditional) x grhaylhdxhybridxentropyxconservsxtoxprims x entropy+hybrid version of grhaylhdxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)

11

GRHayLHDxouterxboundaries (conditional) x grhaylhdxhybridxentropyxouterxboundaries x entropy+hybrid version of grhaylhdxouterxboundaries x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)

GRHayLHDxevaluatexsourcesxrhs (conditional) x grhaylhdxhybridxentropyxevaluatexsourcesxrhs x entropy+hybrid version of grhaylhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs

11

GRHayLHDxevaluatexfluxesxrhs (conditional) x grhaylhdxhybridxentropyxevaluatexfluxesxrhs x entropy+hybrid version of grhaylhdxevaluatexfluxesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 entxstarxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxflux
11 x 11 x  11 entxstarxrhs

11

GRHayLHDxperturbxprimitives (conditional) x grhaylhdxhybridxentropyxperturbxprimitives x entropy+hybrid version of grhaylhdxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities(everywhere)

11

GRHayLHDxperturbxconservatives (conditional) x grhaylhdxhybridxentropyxperturbxconservatives x entropy+hybrid version of grhaylhdxperturbxconservatives

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)

11

GRHayLHDxprimsxtoxconservs (conditional) x grhaylhdxhybridxprimsxtoxconservs x hybrid version of grhaylhdxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 grmhdxvelocities
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

GRHayLHDxconservsxtoxprims (conditional) x grhaylhdxhybridxconservsxtoxprims x hybrid version of grhaylhdxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

GRHayLHDxouterxboundaries (conditional) x grhaylhdxhybridxouterxboundaries x hybrid version of grhaylhdxouterxboundaries

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

GRHayLHDxevaluatexsourcesxrhs (conditional) x grhaylhdxhybridxevaluatexsourcesxrhs x hybrid version of grhaylhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs

11

GRHayLHDxPrim2Con2Prim (conditional) x convertxhydrobasextoxgrhaylhd x convert hydrobase initial data to grhaylhd variables x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::vel
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxvelocities(everywhere)

11

GRHayLHDxevaluatexfluxesxrhs (conditional) x grhaylhdxhybridxevaluatexfluxesxrhs x hybrid version of grhaylhdxevaluatexfluxesxrhs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 grmhdxconservativesxrhs

11

GRHayLHDxperturbxprimitives (conditional) x grhaylhdxhybridxperturbxprimitives x hybrid version of grhaylhdxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)

11

GRHayLHDxperturbxconservatives (conditional) x grhaylhdxhybridxperturbxconservatives x hybrid version of grhaylhdxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)

11

GRHayLHDxprimsxtoxconservs (conditional) x grhaylhdxtabulatedxentropyxprimsxtoxconservs x entropy+tabulated version of grhaylhdxprimsxtoxconservs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

GRHayLHDxconservsxtoxprims (conditional) x grhaylhdxtabulatedxentropyxconservsxtoxprims x entropy+tabulated version of grhaylhdxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

GRHayLHDxouterxboundaries (conditional) x grhaylhdxtabulatedxentropyxouterxboundaries x entropy+tabulated version of grhaylhdxouterxboundaries

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

GRHayLHDxevaluatexsourcesxrhs (conditional) x grhaylhdxtabulatedxentropyxevaluatexsourcesxrhs x entropy+tabulated version of grhaylhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs
11 x 11 x  11 yexstarxrhs

11

GRHayLHDxevaluatexfluxesxrhs (conditional) x grhaylhdxtabulatedxentropyxevaluatexfluxesxrhs x entropy+tabulated version of grhaylhdxevaluatexfluxesxrhs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 entxstarxrhs(interior)
11 x 11 x  11 yexstarxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 entxstarxflux
11 x 11 x  11 yexstarxflux
11 x 11 x  11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs
11 x 11 x  11 yexstarxrhs

11

GRHayLHDxperturbxprimitives (conditional) x grhaylhdxtabulatedxentropyxperturbxprimitives x entropy+tabulated version of grhaylhdxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)

11

GRHayLHDxperturbxconservatives (conditional) x grhaylhdxtabulatedxentropyxperturbxconservatives x entropy+tabulated version of grhaylhdxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)

11

GRHayLHDxPrim2Con2Prim (conditional) x grhaylhdxperturbxprimitives x perturb initial primitive data x

11 x 11 After: 11 convertxhydrobasextoxgrhaylhd
11 x 11 Before: 11 grhaylhdxprimsxtoxconservs
11 x 11 Type: 11 group

11

GRHayLHDxprimsxtoxconservs (conditional) x grhaylhdxtabulatedxprimsxtoxconservs x tabulated version of grhaylhdxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

GRHayLHDxconservsxtoxprims (conditional) x grhaylhdxtabulatedxconservsxtoxprims x tabulated version of grhaylhdxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

GRHayLHDxouterxboundaries (conditional) x grhaylhdxtabulatedxouterxboundaries x tabulated version of grhaylhdxouterxboundaries x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

GRHayLHDxevaluatexsourcesxrhs (conditional) x grhaylhdxtabulatedxevaluatexsourcesxrhs x tabulated version of grhaylhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 yexstarxrhs

11

GRHayLHDxevaluatexfluxesxrhs (conditional) x grhaylhdxtabulatedxevaluatexfluxesxrhs x tabulated version of grhaylhdxevaluatexfluxesxrhs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 yexstarxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 grmhdxconservativesxrhs
11 x 11 x  11 yexstarxflux
11 x 11 x  11 yexstarxrhs

11

GRHayLHDxperturbxprimitives (conditional) x grhaylhdxtabulatedxperturbxprimitives x tabulated version of grhaylhdxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

GRHayLHDxperturbxconservatives (conditional) x grhaylhdxtabulatedxperturbxconservatives x tabulated version of grhaylhdxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)

11

GRHayLHDxPrim2Con2Prim (conditional) x grhaylhdxprimsxtoxconservs x compute conservative variables from primitives x

11 x 11 After: 11 convertxhydrobasextoxgrhaylhd
11 x 11 Type: 11 group

GRHayLHDxPrim2Con2Prim (conditional) x grhaylhdxconservsxtoxprims x compute primitive variables from conservatives x

11 x 11 After: 11 grhaylhdxprimsxtoxconservs
11 x 11 Type: 11 group

11

GRHayLHDxPrim2Con2Prim (conditional) x convertxgrhaylhdxtoxhydrobase x convert grhaylhd-native variables to hydrobase x

11 x 11 After:  11 grhaylhdxconservsxtoxprims
11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::vel(everywhere)
11 x 11 x  11 hydrobase::wxlorentz(everywhere)

11

HydroBasexCon2Prim (conditional) x grhaylhdxcon2prim x compute primitive variables from conservatives x

11 x 11 Type: 11 group

11

GRHayLHDxCon2Prim (conditional) x grhaylhdxemptyxfunction x sync conservative variable ghost zones for c2p routine x

11 x 11 Language: 11 c
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 x  11 entxstar
11 x 11 Type:  11 function

11

Aliased Functions x

11 Alias Name: xxxxxxx  11 Function Name:
11 GRHayLHDxemptyxfunction 11 GRHayLHDxsyncxconservatives